### v1

@@ 4.1 Stochastic Orders
We now define three classes of sets, which we will use to define three stochastic orders. A set $U\subset\Theta$ is an upper set if $\theta\in U$ whenever $\theta\in\Theta$, $\theta\geq\delta$ and $\delta\in U$. A set $U\subset\Theta$ is an upper orthant if it is of the form $\{\,\theta\in\Theta\mid\theta_{1}>a_{1},\theta_{2}>a_{2},\ldots,\theta_{d}>a_{d}\,\}$ for some $a\in\mathbb{R}^{d}$. A set $U\subset\Theta$ is an upper projection if it is of the form $\{\,\theta\in\Theta\mid\theta_{i}>a\,\}$ for some $a\in\mathbb{R}$ and $i\in[d]$. [^11] Upper sets and upper orthants are standard definitions in the literature on stochastic orders, upper projection is not [48]. Note that we could have equivalently defined each of these classes of sets over $\mathbb{R}^{d}$, in which case we would have said that a set $U\subset\Theta$ is an upper set if there exists an upper set $\tilde{U}\in\mathbb{R}^{d}$ such that $U=\Theta\cap\tilde{U}$.
@@
Let $P^{L},P^{H}$ be two probability measures over $\Theta$ and $P^{L}_{i},P^{H}_{i}$ be their respective marginals on the $i$ coordinate. Say:
@@
It is immediate that every upper projection is an upper orthant, but not vice versa, and that every upper orthant is an upper set, but not vice versa. This proves the following result, which states that the orders in Definition 3 are ordered from the strongest, or most restrictive, to the weakest. [^12] This relationship is also noted by [48] who do not formally define coordinatewise polarization but mention that upper orthant polarization implies univariate polarization on each coordinate.
@@
Another way to define stochastic orders is via inequalities of expectations over a class of functions. Given a set of functions $\mathcal{U}\subset\mathbb{R}^{\Theta}$, and a partial order $\succeq$ over $\Delta(\Theta)$, say that $\succeq$ is generated by $\mathcal{U}$ if
$P^{L}\preceq P^{H}\iff E^{L}[u]\leq E^{H}[u]\quad\forall u\in\mathcal{U}$
Here $E^{L}[u]$ and $E^{H}[u]$ are the expectations of $u$ according to the distributions $P^{L}$ and $P^{H}$, respectively. We will use this notation throughout the paper. Table 2 below summarizes the generating functions for the three stochastic orders defined in 3. Here, by increasing function we mean a function $u:\Theta\rightarrow\mathbb{R}$ that is increasing according to the order $\geq$. A product of nonnegative univariate increasing functions is a function $u:\Theta\rightarrow\mathbb{R}$ such that $u(\theta)=\Pi_{i=1}^{d}u_{i}(\theta_{i})$ where each $u_{i}$ is nonnegative and increasing, and a sum of nonnegative univariate increasing functions is defined similarly. The generating functions for the orders $\succeq_{st}$ and $\succeq_{uo}$ are presented in [48], and the generating functions for $\succeq_{cw}$ are derived in Lemma 2 below. [^13] By linearity of the expectation, if an order $\succeq$ is generated by a set $\mathcal{U}$ of functions then it is also generated by the set of all functions that are convex combinations of functions in $\mathcal{U}$.
@@
Stochastic Order Generating Functions
$\preceq_{st}$ Increasing functions
$\preceq_{uo}$ Products of nonnegative
@@
$\preceq_{cw}$ Sums of univariate increasing functions
@@
The order $\preceq_{cw}$ is generated by the set
$\mathcal{U}=\{\,u\in\mathbb{R}^{\Theta}\mid u(\theta)=\sum_{i=1}^{d}u_{i}(\theta_{i})\text{ where }u_{i}:\Theta_{i}\rightarrow\mathbb{R}\text{ are increasing}\,\}$
4.2 Limit Posteriors
Consider an agent with prior $P\in\Delta(\Theta)$ observing an infinite sequence $X_{1},X_{2},\ldots$ of i.i.d. signals. Let $\ell$ be the likelihood function of each of these signals, with $\ell(x;\theta)=\Pr[X=x|\theta]$. [^14] Outside this section we fix the signal realization $x$ and simply write $\ell(\theta)$. Let $\hat{\theta}$ be the true state of the world, the identified set is
Lemma 3.
Lemma 3 implies that the limit belief following an infinite sequence of signals is almost surely equal to the posterior belief that would arise upon observing a single signal that revealed that the state is in the identified set and nothing else. Formally, say that $X$ is a partitional signal if there exists a partition $\mathcal{G}$ of $\Theta$ and a one-to-one mapping $x:\mathcal{G}\rightarrow\mathcal{X}$ such that for every $\theta\in G\in\mathcal{G}$:
$\ell(x(G);\theta)=1\text{ and }\ell(x^{\prime};\theta)=0\text{ for all }x^{\prime}\neq x(G)$
Definition 2’.
Fix a partitional signal $X$ with partition $\mathcal{G}$. Let $P^{L},P^{H}\in\Delta(\Theta)$ be the agents’ priors and let $Q^{L},Q^{H}$ be their respective posteriors upon observing a realization $\Gamma$ of $X$. Say that limit polarization occurs for the identified set $\Gamma$ if
$Q^{L}\prec\ P^{L}\prec P^{H}\prec Q^{H}$
This formulation will be our working definition. One advantage is that it clarifies that limit polarization is a special case of one-shot polarization. An immediate implication is that if limit polarization is possible then so is one-shot polarization. Conversely, if one shot-polarization is impossible then limit polarization is impossible as well.
5 (Im)Possibility of Polarization
This section presents the results about the possibility and impossibility of polarization that are summarized in Table 1. We begin with the coordinatewise order.
@@ 4.2 Limit Posteriors
Limit $\succeq_{cw}$-polarization is possible.
We show in Section 6 that many identified sets can lead to coordinatewise polarization. The simplest, which works in any dimension, is the set $\Gamma=\{\underaccent{\bar}{\theta},\bar{\theta}\}$ that includes just the minimal and maximal states. If agent $L$’s prior significantly favors $\underaccent{\bar}{\theta}$ over $\bar{\theta}$ and is uniform over the remaining states, and agent $H$’s is the mirror image, then this set may lead to coordinatewise polarization. The following example illustrates in a $2\times 2$ state space and the proof of Theorem 1 generalizes to any finite product space $\Theta$.
Example 1.
Let $\Theta=\{x_{1},x_{2}\}\times\{y_{1},y_{2}\}$ with the priors given in Figure 2.
[…]
Figure 2: Priors for Example 1.
Polarization occurs for the identified set $\Gamma=\{(x_{1},y_{1}),(x_{2},y_{2})\}$, because the marginals on any coordinate $i$ satisfy:
$Q^{L}_{i}=(3/4,1/4)\prec_{st}P^{L}_{i}=(5/8,3/8)\prec_{st}P^{H}_{i}=(3/8,5/8)\prec_{st}Q^{H}_{i}=(1/4,3/4)$
We saw above that if limit polarization is possible according to a stochastic order then one-shot polarization according to that order is possible as well. Hence, an immediate corollary of Theorem 1 is that one-shot coordinatewise polarization is possible, thereby completing the first row of Table 1.
We now turn to the upper orthant order. First, consider a $2\times 2$ state space as in Example 1. This state space has four upper orthants, illustrated in Figure 3. Note that the priors and identified set $\Gamma$ presented in Example 1 do not lead to upper orthant polarization, since for both agents the mass on the singleton upper set $U_{4}$ increases after observing the signal. Hence $Q^{L}$ is not upper orthant dominated by her $P^{L}$ so upper orthant polarization does not occur, and this will the be the case for any identified set that includes $\bar{\theta}=(x_{2},y_{2})$. On the other hand, if the identified set does not include $\bar{\theta}$ then for both agents the mass on $U_{4}$ decreases after observing the signal. But this implies that $P^{H}$ is not upper orthant dominated by $Q^{H}$ so that upper orthant polarization does not occur. This argument extends to any state space $\Theta$ and leads to the following result.
@@
Theorem 2.
Limit $\succeq_{uo}$-polarization is impossible.
With partitional signals, agents’ posteriors place more mass than their priors on any state in the identified set, and zero mass outside. Thus, for any singleton event, agents’ beliefs move in the same direction. This was not an issue for coordinatewise polarization—where none of the relevant events (upper projections) are singletons—but it rules out limit upper orthant polarization.
Without the restriction that the signal is partitional, agents’ beliefs may move in opposite directions also over singleton events, and this allows for one-shot upper orthant polarization.
Theorem 3.
One-shot $\succeq_{uo}$-polarization is possible.
The proof builds on the observation that the posterior probability of a state exceeds its prior if and only if its likelihood is larger than the expected likelihood under the prior. Formally, if $P$ is the prior and $Q$ the posterior after observing a realization with likelihood function $\ell$, then for any $\theta$, $Q(\theta)>P(\theta)$ iff $\ell(\theta)>E_{P}[\ell]$. Lemma 9 in the appendix gives a more general version.
To see how this enables upper orthant polarization, consider the $2\times 2$ space, and a signal realization $x$ with the likelihood function $\ell$ presented in Figure 4. If agent $L$’s prior concentrates around the minimal state $\underaccent{\bar}{\theta}$ then her expected likelihood over the state space will be close to $\ell(\underaccent{\bar}{\theta})=1$. Sine this exceeds $\ell(\bar{\theta})=\frac{1}{2}$, her belief on $\bar{\theta}$ decreases after observing $x$. Conversely, we can construct agent $H$’s prior so that her expected likelihood is below $\frac{1}{2}$, leading her belief on $\bar{\theta}$ to increase. Thus, beliefs diverge on the singleton upper orthant. With this likelihood function and suitable priors, beliefs may diverge on all upper orthants.
Note that unlike the symmetric likelihood function that generated coordinatewise polarization, here a key feature is that the likelihoods of $\underaccent{\bar}{\theta}$ and $\bar{\theta}$ differ. This allows for agents to initially update in opposite directions but means that the state is identified so agents’ limit beliefs converge to the true state.
@@
We complete this section by showing that one-shot polarization according to the stochastic dominance order is impossible. An immediate corollary is that limit polarization according to this order is also impossible, thereby completing the last row of Table 2.
@@
Theorem 4 extends Theorem 1 of BHK from finite one-dimensional state spaces to any finite multidimensional state space. [^15] The proof does not rely on the assumption that $\Theta$ is a product space. The proof follows theirs but instead of using induction on the states as in BHK, the induction argument is on the sizes of upper sets from $1$ to $|\Theta|$.
To get an intuition for why one-shot upper orthant polarization is possible while one-shot stochastic polarization is not, recall that all upper orthants are upper sets but not vice-versa. In the $2\times 2$ state space there is one upper set that is not an upper orthant, the “upper L” presented in Figure 5. For polarization to occur the high agent’s beliefs must increase on the upper set $\{\bar{\theta}\}$ and decrease on $\{\underaccent{\bar}{\theta}\}$, which is the complement of the upper $L$. The opposite must occur for the low agent, but this is violates general restriction implied by Bayes rule.

### v2

@@ 5 Constraints on Bayesian Polarization
A set $U\subset\Theta$ is an upper set if $\theta\in U$ whenever $\theta\in\Theta$, $\theta\geq\delta$ and $\delta\in U$. A set $U\subset\Theta$ is an upper orthant if it is of the form $\{\,\theta\in\Theta\mid\theta\gg a\,\}$ for some $a\in\mathbb{R}^{d}$. A set $U\subset\Theta$ is an upper projection if it is of the form $\{\,\theta\in\Theta\mid\theta_{i}>a\,\}$ for some $a\in\mathbb{R}$ and $i\in[d]$. [^17] Upper sets and upper orthants are standard in the stochastic orders literature; upper projections are not. Equivalently, these classes of sets can be defined over $\mathbb{R}^{d}$, in which case we would say that a set $U\subset\Theta$ is an upper set if there exists an upper set $\tilde{U}\subset\mathbb{R}^{d}$ such that $U=\Theta\cap\tilde{U}$.
@@
Let $P^{L},P^{H}$ be two probability measures over $\Theta$ and $P^{L}_{i},P^{H}_{i}$ be their respective marginals on the $i$ coordinate. We say that:
@@
Stochastic dominance and upper orthant dominance are the two standard generalizations of univariate stochastic dominance to multiple dimensions, with the former being the most common [39].
It is immediate that every upper projection is an upper orthant, but not vice versa, and that every upper orthant is an upper set, but not vice versa. This proves the following result, which states that the orders in Definition 7 are ordered from the strongest, or most restrictive, to the weakest. [^18] This relationship is also noted by [39], who do not formally define coordinatewise dominance but mention that upper orthant dominance implies univariate stochastic dominance on each coordinate.
@@
The stochastic orders defined above belong to the class of integral stochastic orders [43, 29]. An order $\succeq$ in this class is characterized by a generator $\mathcal{U}$, which is a set of functions such that $P^{L}\preceq P^{H}$ if and only if $E^{L}[u]\leq E^{H}[u]$ for all $u\in\mathcal{U}$. Here and below, $E^{L}[u]$ and $E^{H}[u]$ are the expectations of $u$ according to the distributions $P^{L}$ and $P^{H}$, respectively. Table 2 summarizes the generators for the three orders. By an increasing function, we mean a function $u:\Theta\rightarrow\mathbb{R}$ that is increasing according to the order $\geq$. A product of nonnegative univariate increasing functions is a function $u:\Theta\rightarrow\mathbb{R}$ such that $u(\theta)=\prod_{i=1}^{d}u_{i}(\theta_{i})$ where each $u_{i}$ is nonnegative and increasing on $\Theta_{i}$, and a sum of univariate increasing functions is defined similarly. The characterizations for $\succeq_{st}$ and $\succeq_{uo}$ are standard [39], while the characterization for $\succeq_{cw}$ is derived in Lemma 4.191919By the linearity of expectation, if an order is generated by $\mathcal{U}$, it is also generated by the convex cone spanned by $\mathcal{U}$.
@@
Stochastic Order Generator
$\succeq_{st}$ Increasing functions
$\succeq_{uo}$ Products of nonnegative
@@
$\succeq_{cw}$ Sums of univariate increasing functions
@@
The order $\succeq_{cw}$ is generated by the set
$\mathcal{U}=\{\,u\in\mathbb{R}^{\Theta}\mid u(\theta)=\sum_{i=1}^{d}u_{i}(\theta_{i})\text{ where }u_{i}:\Theta_{i}\rightarrow\mathbb{R}\text{ are increasing}\,\}.$
We now turn to polarization in the upper orthant order. The proof of Theorem 3 can easily be modified to apply to the closely related lower orthant order.
@@
One-shot $\succeq_{uo}$-polarization is possible. Limit $\succeq_{uo}$-polarization is impossible.
To understand the impossibility result, recall that limit beliefs concentrate on the identified set—the mass on all states in the identified set increases while the mass on all states outside the set vanishes. Consider a $2\times 2$ state space as in Example 1. This state space has four upper orthants, illustrated in Figure 6. Any identified set $\Gamma$ that contains $\bar{\theta}=(x_{2},y_{2})$ cannot lead to limit upper orthant polarization, because for both agents the mass on the singleton upper orthant $U_{4}=\{\bar{\theta}\}$ increases after observing the signal. In this case, $Q^{L}$ is not upper orthant dominated by $P^{L}$, so upper orthant polarization does not occur. On the other hand, if the identified set does not include $\bar{\theta}$, then for both agents the mass on $U_{4}$ decreases. This implies that $P^{H}$ is not upper orthant dominated by $Q^{H}$, again precluding polarization. This argument extends to any state space $\Theta$, establishing the impossibility of limit upper orthant polarization.
@@
With partitional signals, agents must update in the same direction on any singleton event. By contrast, general signals allow beliefs to move in opposite directions even on singletons. We show that this allows for one-shot upper orthant polarization. The proof builds on the observation that an agent’s belief on a state increases if and only if the state’s likelihood exceeds the expected likelihood according to the agent’s prior. This is a special case of the following simple result.
Lemma 5.
Let $P\in\Delta(\Theta)$ be a prior belief with full support over $\Theta$ and $Q\in\Delta(\Theta)$ be the corresponding posterior after observing a realization $x$ of a signal $X$ with likelihood $\ell(\theta)=\Pr[X=x\mid\theta]$. Then, for every nonempty $A\subset\Theta$: $Q(A)>P(A)$ if and only if $E_{P}[\ell|A]>E_{P}[\ell]$.
To see why one-shot upper orthant polarization is possible, consider the $2\times 2$ space and a signal realization $x$ with the likelihood function $\ell$ presented in Figure 7. If $L$’s prior mostly concentrates on $\underaccent{\bar}{\theta}$ (where $\ell=1$), their expected likelihood is close to $1$, causing their belief on $\bar{\theta}$ (where $\ell=1/2$) to fall after observing $x$. Conversely, we can construct $H$’s prior so that their expected likelihood is below $1/2$, so that their belief on $\bar{\theta}$ increases after observing $x$. With suitable priors, this divergence on the singleton upper orthant extends to all upper orthants. Note that unlike the symmetric signals that generate limit coordinatewise polarization, here the likelihoods of $\underaccent{\bar}{\theta}$ and $\bar{\theta}$ differ; this asymmetry enables one-shot polarization but also identifies the state. Thus, if the agents observe an infinite sequence of i.i.d. signals, their beliefs will eventually concentrate on the true state.
@@
Since the upper orthant order is weaker than stochastic dominance, the impossibility of limit upper orthant polarization implies that limit stochastic dominance polarization is impossible as well. The following result then completes Table 1.
@@
Theorem 4 extends Theorem 1 of BHK from finite one-dimensional state spaces to arbitrary finite multidimensional state spaces. [^20] The proof does not rely on the assumption that $\Theta$ is a product space. We first briefly review their argument. Consider a finite one-dimensional state space $\theta_{1}<\theta_{2}<\ldots<\theta_{n}$. For polarization in stochastic dominance to occur, the low agent’s beliefs must increase on the lowest state while the high agent’s beliefs decrease, and the opposite must occur on the highest state. [^21] That is, $Q^{L}(\theta_{1})\geq P^{L}(\theta_{1})$, $Q^{H}(\theta_{1})\leq P^{H}(\theta_{1})$, while $Q^{L}(\theta_{n})\leq P^{L}(\theta_{n})$, $Q^{H}(\theta_{n})\geq P^{H}(\theta_{n})$. BHK show that if these occur simultaneously, then beliefs at these states do not move at all (we extend this argument in Proposition 1 below). They then proceed by induction to show that beliefs are constant on every state—i.e., $Q^{L}(\theta_{i})=P^{L}(\theta_{i})$ and $P^{H}(\theta_{i})=Q^{H}(\theta_{i})$ for every $i\in[n]$—so polarization does not occur.
Our proof adapts this strategy to multidimensional state spaces. Instead of inducting on states, we induct on the size of upper sets. That is, we show that for every upper set $U$, $Q^{H}(U)=P^{H}(U)$ and $Q^{L}(U)=P^{L}(U)$ by induction on the size of $U$. The key property is that if $U$ is an upper set of size $m$, then removing any of its minimal elements leads to an upper set of size m1m-1.
To understand why one-shot upper orthant polarization is possible while one-shot stochastic polarization is not, consider again the $2\times 2$ state space. Here, there is one upper set that is not an upper orthant: the “upper L” presented in Figure 8 (recall that all upper orthants are upper sets, but not vice versa). For $\succeq_{st}$-polarization to occur, the high agent’s beliefs must increase on the upper set $\{\bar{\theta}\}$ and decrease on $\{\underaccent{\bar}{\theta}\}$, which is the complement of the upper L. The opposite must occur for the low agent. However, this pattern of divergence violates a general restriction implied by Bayesian updating, which we formalize in the following proposition.